Q5 Consider the following data on Malaysia's GDP :
Year Norminal GDP (in billion) GDP deflator
2005 495,239 189
2006 546,343 197
a) what is the growth rate of norminal GDP between 2005 and 2006 ( note: the growth rate is the percentage change from one period to the next )
b) what is the growth rate of the GDP deflator between 2005 and 2006?
c)what was real GDP in 2005 measured in 1987 prices?
d) what was real GDP in 2006 measured in 1987 prices?
e)what was the growth rate of real GDP between 2005 and 2006?
f)was the growth rate of nominal GDP higher or lower than the growth rate of real GDP? Explain
1
Expert's answer
2020-12-08T10:16:38-0500


a) growth rate = (nominal GDP in 2006 - nominal GDP in 2005)"\/" nominal GDP in 2005 "\times 100"

"=\frac{546343-495239}{495239}\times 100=10.32\%"

b) growth rate of GDP deflator "=\frac{197-189}{189}\times 100=4.23\%"

c) GDP deflator, 2005 "=" nominal GDP of 2005 "\/" real GDP of 2005 "\times" "100"

"189=\frac{495239}{real GDP_{2005}}\times 100"

real "GDP_{2005} =\frac{495239}{189}\times 100= 262031.22"


d) real "GDP_{2006} =\frac{546343}{197}\times 100=277331.47"


e) growth rate of real "GDP =\frac{277331.47-262031.22}{262031.22}\times100=5.84\%"

f) the growth rate of nominal GDP was higher, since nominal "GDP(10.32\%" ) is greater than real "GDP(5.84\%)"